OpenStudy (anonymous):

Please help crashing for finals! The length of the rectangle is 4t+5 and its height is t^6, where t is time in seconds and the dimensions are in inches. Find the rate of change of area, A, with respect to time.

OpenStudy (anonymous):

A= l*h = (4t+5)(t^6) dA/dt = d/dt (4t^7 +5t^6 )

OpenStudy (anonymous):

my teacher gave me the answer it is dA/dt=t^3(30+28t) but im not so sure as to how to do the problem

OpenStudy (anonymous):

aka 28t^6 +30t^5

OpenStudy (anonymous):

thats what i got when i first did the problem what you gave me just now

OpenStudy (anonymous):

dA/dt=t^5(30+28t) think you have a typo there..

OpenStudy (anonymous):

or in the original problem...

OpenStudy (anonymous):

yeah you know what i was thinking that too. so doing the problem just as i had done it taking the derivative of the problem thats how we go about these problems right? simply taking the derivative??

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

:( no i need a definite answer because i have a final in a couple of days and i NEED to get a good score on it

OpenStudy (anonymous):

"so doing the problem just as i had done it taking the derivative of the problem thats how we go about these problems right?" I have no idea what that word salad might mean... so I decline to comment on it aside from a provisional 'sure'

OpenStudy (anonymous):

okay then thanks